# The Bayesian Separation Principle for Data-driven Control

### v1

@@ V-C Explicit expressions of MPC cost
Linear case. Consider now systems that are linear in the past inputs and outputs as in (22). Let $Y$ denote the vector with the future outputs $y(t),\dots,y(t+T-1)$ which we would like to drive to $y_{\text{ref}}(t),\dots,y_{\text{ref}}(t+T-1)$. For our purposes, using the formulas in [29, Sec. 3.2], it is important to write the measurements model as follows
$Y=A(\theta)U+B(\theta)Y_{-}+C(\theta)U_{-}+D(\theta)E$
@@
where the entries of matrices $A$, $B$, $C$, $D$ are nonlinear functions of $\theta$ such that the $k$th row of $A(\theta)U+B(\theta)Y_{-}+C(\theta)U_{-}$ is the $k$-step ahead predictor of $y(t+k-1)$. In addition, the vectors $U_{-}$, $Y_{-}$ contain the past inputs $u(t-1),u(t-2),\dots$, and outputs $y(t-1),y(t-2),\dots$ while $U\in\mathbb{R}^{T}$ is the vector containing the inputs to be optimized, namely $u(t),\dots,u(t+T-1)$. Conditional on the identification data $\mathcal{D}$, the model becomes
@@
$\mathbb{E}[\|Y_{\text{ref}}-A(\theta^{P})U-B(\theta^{P})Y_{-}-C(\theta^{P})U_{-}\|^{2}]$
$+\mathbb{E}[\|D(\theta^{P})E\|^{2}]$
$+2\mathbb{E}\Big{[}\sum_{i}[A(\theta^{P})U+B(\theta^{P})Y_{-}+C(\theta^{P})U_{-}]_{i}[D(\theta^{P})E]_{i}\Big{]}$
where $Y_{\text{ref}}\in\mathbb{R}^{T}$ contains the reference outputs $y_{\text{ref}}(t),\dots,y_{\text{ref}}(t+T-1)$. Note that the second term on the rhs does not depend on $U$ and the last term is zero since the zero-mean noise $E$ is independent of $\theta^{P}$. This, combined with Remark 3, permits to write the conditional MPC cost as
$J_{u}^{P}$ $=$ $\|Y_{\text{ref}}-A(\theta^{P})U-B(\theta^{P})Y_{-}-C(\theta^{P})U_{-}\|_{Q}^{2}$ (38)
$+$ $\|U_{\text{ref}}-U\|_{R}^{2}$
where $Q:=\mathrm{blkdiag}\{Q_{\ell}\}_{\ell=0,\dots,T-1}$, $R:=\mathrm{blkdiag}\{R_{\ell}\}_{\ell=0,\dots,T-1}$ while $U_{\text{ref}}$ contains the reference inputs $u_{\text{ref}}(t),\dots,u_{\text{ref}}(t+T-1)$. The posterior mean of (38) is
$\hat{J}_{u}$ $=U^{\top}\mathbb{E}[A^{\top}QA\,|\,\mathcal{D}]U$
$-2U^{\top}\!(\mathbb{E}[A^{\top}|\,\mathcal{D}]QY_{\text{ref}}-\mathbb{E}[A^{\top}\!QB\,|\,\mathcal{D}]Y_{-}\!$
@@
Even if the entries of matrices $A$, $B$, $C$ are nonlinear functions of the random vector $\theta^{P}$, the terms $\mathbb{E}[A^{\top}QA\,|\,\mathcal{D}]$, $\mathbb{E}[A^{\top}|\,\mathcal{D}]$, $\mathbb{E}[A^{\top}\!C\,|\,\mathcal{D}]$ appearing in $U^{*}$ can be computed offline either in closed-form or numerically via a simple Monte Carlo strategy.
VI The Bayesian separation principle in action
@@
VI-A Closed-form MPC cost in a simple linear setting
Let us still assume a scalar output $y(t)\in\mathbb{R}$ with the reference trajectory $Y_{\text{ref}}\in\mathbb{R}^{T}$. Following the notation of the previous section the predictor of $Y_{\text{ref}}$ is
$AU+BY_{-}+CU_{-}$
where $A\in\mathbb{R}^{T\times T}$, $B\in\mathbb{R}^{T\times m}$, $C\in\mathbb{R}^{T\times m-1}$ and $m$ is the number of past inputs and past outpts the dynamic system $f$ depends on. We let $t=1$ so that the vector $U$ contains the optimization variables
$U=\left[\begin{array}[]{c}u_{1}\\ u_{2}\\ \vdots\\ u_{T}\end{array}\right].$
Consider the simplest case $m=1$ with the predictor impulse response coefficients contained in the two-dimensional vector $\theta$. Specifically, the one-step ahead predictor is
$\hat{y}_{1}=\theta_{1}u_{1}+\theta_{2}y_{0},$
@@ VI-A Closed-form MPC cost in a simple linear setting
$\hat{y}_{2}=\theta_{1}u_{2}+\theta_{2}\theta_{1}u_{1}+\theta_{2}^{2}y_{0}$
and so on. One then has
@@
Let us set $T=2$ to further simplify the example. It follows that
@@
After exploiting the identification data $\mathcal{D}$, the vector $\theta$ is Gaussian with mean $\mu$ and covariance $\Sigma$ with $(i,j)$ entry equal to $\sigma_{ij}$ outside the diagonals and to $\sigma_{i}^{2}$ along the diagonal. After some simple calculations based on non central Gaussian moments, one obtains
$\mathbb{E}[A^{\top}|\mathcal{D}]=\left[\begin{array}[]{cc}\mu_{1}&\mu_{1}\mu_{2}+\sigma_{12}\\ 0&\mu_{1}\end{array}\right]$
@@
VI-B Numerical experiment
Figure 1: Left: boxplot of 10000 costs returned by MPC using only the nominal model or exploiting the Bayesian separation principle to compute the input which minimizes the minimum variance estimate of the cost. Center: boxplot of the 10000 costs related to the part of the objective containing the one-step ahead prediction error. Right: boxplot of the 10000 costs related to the part of the objective containing the two-step ahead prediction error.
We are now in a position to exploit our explicit formulas in a numerical example. Let
@@ VI-B Numerical experiment
and $Y_{-}=1$. The MPC cost (IV-B) is defined setting $Q$ to the identity matrix and $R$ to the null matrix (no penalty on the control input is given). Let us compare three different control schemes:
  • Oracle: the realization of $\theta$ is known and used to build the optimal one- and two-step ahead predictors. This defines the optimal target as $J_{u}=0$. In fact, $A$ is full-rank with probability one, hence one can find $U$ that interpolates perfectly $Y_{\text{ref}}$ since no input penalty is specified;
  • Nominal: only the system estimate $\mu$ is used. So, the uncertainty around the predictor is neglected: the closed-form formulas obtained in this section are plugged in (V-C) but setting $\Sigma$ to the null matrix;
  • BSP: the Bayesian separation principle is used. The closed-form formulas reported above are used in (V-C), hence obtaining the input which minimizes the minimum variance estimate of the MPC cost.
A Monte Carlo of 10000 runs is performed. At any run the system is obtained by drawing $\theta$ from the Gaussian distribution of mean $\mu$ and covariance $\Sigma$. Then, the control inputs described above are computed, applied to the true system and the related MPC cost is evaluated. So, we obtain 10000 MPC costs from Nominal and 10000 from BSP. For illustration purposes it is useful also to decompose the MPC objective into two parts $J^{1}_{u}$ and $J^{2}_{u}$ concerning the one-step ahead and the two-step ahead error prediction:
@@
Fig. 1 (left panel) reports the boxplots of the MPC cost $J_{u}$. The average cost from BSP is around 217, almost four times smaller than that returned by Nominal which is around 935. To better understand what is happening, the bottom panels show the boxplots of the two costs $J^{1}_{u}$ and $J^{2}_{u}$. From the center panel, it is interesting to note that Nominal outperforms BSP in one-step-ahead prediction. But the right panel shows the effect of the Bayesian separation principle: it sacrificed one-step performance to improve accuracy over two steps by exploting information about the uncertainty in the predictor parameters. As a result, the overall error is reduced.
VII Conclusion

### v2

@@ 5.2.2 Linear case
Let $Y$ denote the vector with the future outputs $y_{\tau},\dots,y_{\tau+T-1}$ which we would like to drive to $y^{\text{ref}}_{\tau},\dots,y^{\text{ref}}_{\tau+T-1}$. For our purposes, using the formulas in [23, Sec. 3.2], it is important to write the measurements model as follows
$Y=A(\theta)U+B(\theta)Y_{-}+C(\theta)U_{-}+D(\theta)E$ (43)
@@
where the entries of matrices $A$, $B$, $C$, $D$ are nonlinear functions of $\theta$ such that the $k$th row of $A(\theta)U+B(\theta)Y_{-}+C(\theta)U_{-}$ is the $k$-step ahead predictor of $y_{\tau+k-1}$. In addition, the vectors $U_{-}$, $Y_{-}$ contain the past inputs $u_{\tau-1},u_{\tau-2},\dots$, and outputs $y_{\tau-1},y_{\tau-2},\dots$ while $U\in\mathbb{R}^{T}$ is the vector containing the inputs to be optimized, namely $u_{\tau},\dots,u_{\tau+T-1}$. Conditional on the identification data $\mathcal{D}$, the model becomes
@@
$\mathbb{E}[\|Y_{\text{ref}}-A(\theta^{P})U-B(\theta^{P})Y_{-}-C(\theta^{P})U_{-}\|^{2}]+\mathbb{E}[\|D(\theta^{P})E\|^{2}]$
$+2\mathbb{E}\Big[\sum_{i}[A(\theta^{P})U+B(\theta^{P})Y_{-}+C(\theta^{P})U_{-}]_{i}[D(\theta^{P})E]_{i}\Big]$ (44)
where $Y_{\text{ref}}\in\mathbb{R}^{T}$ contains the reference outputs $y^{\text{ref}}_{\tau},\dots,y^{\text{ref}}_{\tau+T-1}$. Note that the second term on the right-hand side of (5.2.2) does not depend on $U$ and the last term is zero since the zero-mean noise $E$ is independent of $\theta^{P}$. This, combined with Remark 3, allows us to write the conditional MPC cost as
$J_{u}^{P}=\|Y_{\text{ref}}-A(\theta^{P})U-B(\theta^{P})Y_{-}-C(\theta^{P})U_{-}\|_{Q}^{2}+\|U_{\text{ref}}-U\|_{R}^{2}$ (45)
where $Q:=\mathrm{blkdiag}\{Q_{\ell}\}_{\ell=0,\dots,T-1}$, $R:=\mathrm{blkdiag}\{R_{\ell}\}_{\ell=0,\dots,T-1}$ while $U_{\text{ref}}$ contains the reference inputs $u^{\text{ref}}_{\tau},\dots,u^{\text{ref}}_{\tau+T-1}$. The posterior mean of (45) is
$\hat{J}_{u}$ $=U^{\top}\mathbb{E}[A^{\top}QA\,|\,\mathcal{D}]U-2U^{\top}\!(\mathbb{E}[A^{\top}|\,\mathcal{D}]QY_{\text{ref}}-\mathbb{E}[A^{\top}\!QB\,|\,\mathcal{D}]Y_{-}\!$
@@
Even if the entries of matrices $A$, $B$, $C$ are nonlinear functions of the random vector $\theta^{P}$, the terms $\mathbb{E}[A^{\top}QA\,|\,\mathcal{D}]$, $\mathbb{E}[A^{\top}|\,\mathcal{D}]$, $\mathbb{E}[A^{\top}\!C\,|\,\mathcal{D}]$ appearing in $U^{*}$ can be computed offline either in closed-form or numerically via a simple Monte Carlo strategy. In particular, such Monte Carlo strategy involves generating a set of samples of $\theta^{P}$ based on its distribution described by (27) and then averaging the evaluated terms across the samples to approximate their expected values.
6 Bayesian separation principle in action: linear setting
@@
6.1 Closed-form MPC cost in a simple linear setting
Consider a scalar output $y(t)\in\mathbb{R}$ and the reference trajectory $Y_{\text{ref}}\in\mathbb{R}^{T}$. Following the notation of the previous section, the predictor of $Y_{\text{ref}}$ is
$A(\theta)U+B(\theta)Y_{-}+C(\theta)U_{-}$
where $A\in\mathbb{R}^{T\times T}$, $B\in\mathbb{R}^{T\times m}$, $C\in\mathbb{R}^{T\times m-1}$ and the vector $U$ contains the $T$ optimization variables. Consider the simplest case where the system memory is just $m=1$, which implies $\theta\in\mathbb{R}^{2}$. The one-step ahead predictor is
$\hat{y}_{1}=\theta_{1}u_{1}+\theta_{2}y_{0},$ (47)
@@ 6.1 Closed-form MPC cost in a simple linear setting
$\hat{y}_{2}=\theta_{1}u_{2}+\theta_{2}\theta_{1}u_{1}+\theta_{2}^{2}y_{0}$ (48)
and so on. More in general, one has
@@
To further simplify the example let $T=2$, hence only the predictors (47) and (48) will enter into the MPC cost. One has
@@
After seeing the identification data $\mathcal{D}$, we assume that the vector $\theta$ is Gaussian with mean $\mu$ with components $\mu_{1},\mu_{2}$. Its covariance $\Sigma$ has $(i,j)$ entry denoted by $\sigma_{ij}$ outside the diagonals and by $\sigma_{i}^{2}$ along the diagonal. After simple calculations based on non-central Gaussian moments, one then obtains
$\mathbb{E}[A^{\top}|\mathcal{D}]$ $=\left[\begin{array}[]{cc}\mu_{1}&\mu_{1}\mu_{2}+\sigma_{12}\\ 0&\mu_{1}\end{array}\right],$
where
@@
6.2 Numerical experiment
Figure 1: Left: boxplot of 10000 costs returned by MPC using only the nominal model (Nominal), exploiting the Bayesian separation principle to compute the input which minimizes the minimum variance estimate of the cost (BSP) or introducing a special (stochastic) weight matrix (W). Center: boxplot of the 10000 costs related to the part of the objective containing the one-step ahead prediction error. Right: boxplot of the 10000 costs related to the part of the objective containing the two-step ahead prediction error.
Let
@@ 6.2 Numerical experiment
and $Y_{-}=1$. The MPC cost (4.2) is defined setting $Q$ to the identity matrix and $R$ to the null matrix, hence no penalty on the control input is assigned. A Monte Carlo of 10000 runs is performed. At any run the system is obtained by drawing $\theta$ from the Gaussian distribution of mean $\mu$ and covariance $\Sigma$. Let us consider three different control schemes:
  • 1.
    Oracle: the true system, i.e. the realization of $\theta$, is known and used to build the optimal one- and two-step ahead predictors. This defines the optimal target as $J_{u}=0$. In fact, $A$ is full-rank with probability one, hence one can find $U$ that interpolates perfectly $Y_{\text{ref}}$;
  • 2.
    Nominal: only the system estimate $\mu$ is used. So, the uncertainty around the predictor is neglected: the closed-form formulas obtained in this section are plugged in (5.2.2) with $\Sigma$ set to the null matrix. The control input is the same at any run, given by
    $\hat{U}^{N}=\left[\begin{array}[]{c}0.5\\ -4\end{array}\right];$ (49)
  • 3.
    BSP: the Bayesian separation principle is used. The closed-form formulas reported above are used in (5.2.2), hence obtaining the input which minimizes the minimum variance estimate of the MPC cost. The control input is now given by
    $\hat{U}^{BSP}\approx\left[\begin{array}[]{c}-0.65\\ 1.48\end{array}\right].$ (50)
At any run, the controls (49) and (50) are applied to the true system and the related MPC cost is evaluated. One thus obtains 10000 MPC costs from Nominal and 10000 from BSP. For illustration purposes it is useful also to decompose the MPC objective into two parts $J^{1}_{u}$ and $J^{2}_{u}$ which describe the one- and two-step ahead error prediction:
@@
Fig. 1 (left panel) reports the boxplots of the MPC cost $J_{u}$. The average cost from BSP is around 217, almost four times smaller than that returned by Nominal which is around 935 (results from a third control strategy denoted by $\bf{W}$} will be described later on). To better understand what is happening, boxplots of the two costs $J^{1}_{u}$ and $J^{2}_{u}$ are also reported. From the center panel, it is interesting to note that Nominal outperforms BSP in one-step-ahead prediction. But the right panel shows the effect of the Bayesian separation principle: it sacrificed one-step performance to improve accuracy over two steps exploiting information on predictor parameters uncertainty. As a result, the overall error is reduced.
6.3 Comparison with the approach in [9]
Working in the linear setting, Theorem 1 in [9, Section 4] reports a closed-form expression of the MPC cost obtained using a special (stochastic) weight matrix $W$ depending on $\theta$. For known $\theta$, such matrix is such that $\sigma^{2}W^{-1}W^{-\top}$ is the covariance of the prediction errors. From (43), the following correspondence is then immediately obtained
$D(\theta)=W^{-1}(\theta),$
and the objective minimized in [9] corresponds to the expectation of the following variation of (45):
$J_{u}^{W}$ $=\Big\|D^{-1}(\theta)\Big(Y_{\text{ref}}-A(\theta^{P})U-B(\theta^{P})Y_{-}-C(\theta^{P})U_{-}\Big)\Big\|_{Q}^{2}$
$\ \ \ \ +\|U_{\text{ref}}-U\|_{R}^{2}$
$=\Big\|\Big(Y_{\text{ref}}-A(\theta^{P})U-B(\theta^{P})Y_{-}-C(\theta^{P})U_{-}\Big)\Big\|_{W^{\top}(\theta)QW(\theta)}^{2}$
$\ \ \ \ +\|U_{\text{ref}}-U\|_{R}^{2}.$ (51)
In our simple example one has
$D(\theta)=\left[\begin{array}[]{cc}1&0\\ \theta_{2}&1\end{array}\right],\quad D^{-1}(\theta)=\left[\begin{array}[]{cc}1&0\\ -\theta_{2}&1\end{array}\right],$
and with $Q=I$ and $R=0$ the cost (51) becomes
$J_{u}^{W}=\left\|\left[\begin{array}[]{cc}1&0\\ -\theta_{2}&1\end{array}\right]Y_{\text{ref}}-U\theta_{1}-\left[\begin{array}[]{c}\theta_{2}\\ 0\end{array}\right]Y_{-}\right\|^{2}$ (52)
which, apart from terms independent of $U$, is equal to
$U^{\top}\left[\begin{array}[]{cc}\theta_{1}^{2}&0\\ 0&\theta_{1}^{2}\end{array}\right]U+2U^{\top}\left(\left[\begin{array}[]{cc}-\theta_{1}&0\\ \theta_{1}\theta_{2}&-\theta_{1}\end{array}\right]Y_{\text{ref}}+\left[\begin{array}[]{c}\theta_{1}\theta_{2}\\ 0\end{array}\right]Y_{-}\right).$
The expected cost $\mathbb{E}[J_{u}^{W}]$ is calculated using
$\mathbb{E}[\theta_{1}|\mathcal{D}]=10,\quad\mathbb{E}[\theta_{2}^{2}|\mathcal{D}]=104,\quad\mathbb{E}[\theta_{1}\theta_{2}|\mathcal{D}]=50.9,$
and its minimizer defines the following control input
$\hat{U}=\frac{1}{104}\left[\begin{array}[]{c}49.1\\ -409\end{array}\right]\approx\left[\begin{array}[]{c}0.47\\ -3.93\end{array}\right].$ (53)
It is very close to the input (49) obtained using the nominal model. The outcomes are indeed very similar as visible from the boxplots of the 10000 MPC costs, labelled W in Fig. 1, obtained using (53) at any Monte Carlo run. One can conclude that the price of using the result in Theorem 1 in [9, Section 4] is to include a stochastic weight matrix that favours small prediction horizons. In fact, as already said, once fixed $\theta$ the matrix $W^{\top}W$ is (proportional to) the inverse of the covariance of the prediction errors whose diagonal elements increase with the prediction horizon. In our example, one indeed has
$\mathbb{E}[W^{\top}W|\mathcal{D}]\approx\left[\begin{array}[]{cc}29.48&-4.95\\ -4.95&1\end{array}\right].$
One could certainly use the matrix $Q$ (here placed at identity) to try to counterbalance this effect but its choice is not so obvious due to the random nature of $W$. Using our control strategy (5.2.2) the effect is instead always transparent, e.g. $Q$ set to the identity always weights the prediction horizons equally.
7 Bayesian separation principle in action: nonlinear setting
In this example, we show how the use of the uncertainty regularizer cost $J_{2}$, as discussed at the end of Section 5.2.1, enhances the performance of the MPC controller with respect to the standard MPC, in which the cost consists only of the “certainty equivalence” term $J_{1}$. We compare the performance of the two controllers in two different numerical experiments.
Consider the controlled Van der Pol oscillator described by the dynamic equations
$\dot{x}_{1}=2x_{2},$
$\dot{x}_{2}=-0.8x_{1}+2x_{2}-10x_{1}^{2}x_{2}+u,$
$y=x_{2},$
where $u$ is the control signal and $y$ is the measured output. The models used in the two experiments are trained with one dataset each, denoted as $\mathcal{D}_{1}$ for the first experiment and $\mathcal{D}_{2}$ for the second experiment. To generate the datasets, in both the experiments we discretize the dynamics using the Runge-Kutta four method, with discretization time $T_{s}=0.05s$, and simulate $100$ trajectories over a time horizon of $T=3s$ per trajectory. The control signal is a random signal with uniform distribution over [1,1][-1,1] for the dataset $\mathcal{D}_{1}$ and $[0,1]$ for the dataset $\mathcal{D}_{2}$. The models are estimated using the Gaussian Process Regression method over the model class of NARX models with memory of $m=4$ past outputs and inputs. To compute the regularized and standard MPC controllers we use the regularized cost $\hat{J}_{u}=J_{1}+J_{2}$ and $\hat{J}_{u}=J_{1}$, where $J_{1}$ and $J_{2}$ are defined in (41) and (42), respectively. We set $Q_{\ell}=1$ and $R_{\ell}=0$ for all the values of the index $\ell$, and we consider a receding horizon of $T=4$ samples. We consider, as output reference, the constant signal $y_{\text{ref}}(t)=0$ for all $t$, i.e., we want to steer the system to the origin.
The MPC problem for NARX models discussed in subsection 5.2.1 can be recast as a nonlinear MPC problem, given the knowledge of the predictor $\mathbb{E}[f(\cdot)|\mathcal{D}]$ and the variance $\mathrm{Var}[f(\cdot)|\mathcal{D}]$, as described next. Let $x_{\tau}=\left[y^{\tau-1}\right]^{T}$,
$x_{\tau+\ell}=\left[\begin{array}[]{c}\mathbb{E}[f(\tilde{z}_{\ell-1}^{*})|\mathcal{D}]\\ \vdots\\ \mathbb{E}[f(\tilde{z}_{0}^{*})|\mathcal{D}]\\ y_{\tau-1}\\ \vdots\\ y_{\tau-m+\ell}\end{array}\right],\,\,\,\ell=1,\ldots,m-1,$
and
$x_{\tau+m}=\left[\begin{array}[]{c}\mathbb{E}[f(\tilde{z}_{m-1}^{*})|\mathcal{D}]\\ \vdots\\ \mathbb{E}[f(\tilde{z}_{0}^{*})|\mathcal{D}]\end{array}\right].$
Then the dynamics for the state $x$ can be written as
$x_{\tau+\ell+1}$ $=\left[\begin{array}[]{ccccc}0&0&&&\\ 1&\ddots&\ddots&&\\ &\ddots&&&\\ &&&&0\\ &&&1&0\end{array}\right]\,x_{\tau+\ell}\,+\,\left[\begin{array}[]{c}\mathbb{E}[f(\tilde{z}_{\ell}^{*})|\mathcal{D}]\\ \vdots\\ \\ \vdots\\ 0\end{array}\right]$ (64)
The problem we aim to solve is
$\underset{u_{\tau},\ldots,u_{\tau+T-1}}{\mathrm{argmin}}\,\,\,\hat{J}_{u}$
subject to(64)
where $\hat{J}_{u}$ is as in (40). This problem can be solved by several numerical toolboxes.
The results of the two experiments are shown in Figure 2. Note that the results show important performance improvements when using the regularizer term $J_{2}$. Indeed, in the first experiment, the model is trained over control signals in the interval [1,1][-1,1]. Approaching the time instant $3.5s$, the unforced system would show strong oscillations, and the output $y$ would go far away from the origin. To compensate for this behavior, the non-regularized MPC estimates an optimal control signal with a large magnitude, outside [1,1][-1,1]. However, these estimates are very unreliable, as the system is not trained with inputs in this interval, causing the poor performance shown in Figure 2, top. The regularized MPC instead, thanks to the regularizer term $J_{2}$, prefers controls in [1,1][-1,1], where the variance of the prediction is small. Although these may not be the optimal choices of the controls in the ideal case where the true system is known, they guarantee that the estimated model gives good predictions and, therefore, that the performance of the system does not degenerate.
In addition, in the second experiment, a similar behavior is shown. The model is trained with inputs in the interval $[0,1]$. The non-regularized MPC, as before, estimates optimal controls outside $[0,1]$, causing the poor performance shown in Figure 2, bottom. The regularized MPC, thanks to the regularizer term $J_{2}$, keeps the controls in the interval $[0,1]$. However, as discussed before, this is not in general the optimal choice for the true system. In particular, this restricts the choices of controls with respect to the first experiment, leading to the worse behavior shown to the right of Figure 2, bottom.
Figure 2: Plots of the controlled $x_{2}$ trajectory using an MPC controller computed using the estimated model. Top: Results of the first experiment, with the model trained using controls in the interval [1,1][-1,1]. The plot using the standard MPC is in blue, the plot using the MPC with regularized cost is in red. Bottom: Results of the second experiment, with the model trained using controls in the interval $[0,1]$. The plot using the standard MPC is in blue, the plot using the MPC with regularized cost is in red.
8 Conclusion